OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (8, 3); m = 6

OpenStudy (anonymous):

what's the point slope form? \[y-y_1=m(x-x_1)?\]

OpenStudy (dumbcow):

lol

OpenStudy (anonymous):

Yes it is apparently: so you have your point (8;3), so x1=8 and y1=3, and m=6. Thus: \[y-3=6(x-8)\]

OpenStudy (anonymous):

@dumbcow why lol?

OpenStudy (anonymous):

: Suppose y varies directly with x, and y = 8 when x = –6. What direct variation equation relates x and y? What is the value of y when x = –2?

OpenStudy (anonymous):

Get it in y=mx+b form first.

OpenStudy (anonymous):

or is this a diff question?

OpenStudy (anonymous):

y=8+-6x

OpenStudy (anonymous):

" Suppose y varies directly with x, and y = 8 when x = –6. What direct variation equation relates x and y? What is the value of y when x = –2?" Is this a separate question from your original one?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

.....

OpenStudy (anonymous):

I had to look this up.

OpenStudy (anonymous):

http://www.freemathhelp.com/direct-variation.html

OpenStudy (anonymous):

-6/8 = -2/y This is a direct variation. Cross multiply and solve for y.

OpenStudy (anonymous):

-6y = -16

OpenStudy (anonymous):

its good explaining 0/375

OpenStudy (anonymous):

0.375

OpenStudy (anonymous):

Nikita, would you prefer just a numerical answer? That will help so much more for your exams.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Believe me none of us remembers every single little thing from algebra. Lol I had to google direct variation of a line.

OpenStudy (anonymous):

Suppose y varies directly with x, and y = 8 when x = –6. What direct variation equation relates x and y? What is the value of y when x = –2? A. y = –0.75x; 1.50 B. y = –1.33x; 2.67 C. y = 1.33x; –2.67 D. y = 0.13x; –0.25

OpenStudy (anonymous):

Hmm... still thinking

OpenStudy (anonymous):

y = k*x Solve for k. 8 = k *(-6) k = -8/6 = -4/3 y = (-4/3)*x Plug in x=-2 into x and solve. y=(-4/3)*(-2) y=8/3 I dunno, I think this is wrong.

OpenStudy (anonymous):

b/c y=2.67 this way :-/

OpenStudy (anonymous):

Just pick A. B/c the method is absolutely correct. Someone *'ed up on writing the answer choices.